# Universal Vessel Segmentation for Multi-Modality Retinal Images

### v3

@@ Abstract
We identify two major limitations in the existing studies on retinal vessel segmentation: (1) Most existing works are restricted to one modality, i.e, the Color Fundus (CF). However, multi-modality retinal images are used every day in the study of retina and diagnosis of retinal diseases, and the study of vessel segmentation on the other modalities is scarce; (2) Even though a few works extended their experiments to limited new modalities such as the Multi-Color Scanning Laser Ophthalmoscopy (MC), these works still require finetuning a separate model for the new modality. The finetuning will require extra training data, which is difficult to acquire. In this work, we present a novel universal vessel segmentation model (UVSM) for multi-modality retinal images. Not only do we perform the study on a much wider range of modalities, but we also propose a universal model to segment the vessels in all these commonly-used modalities. Despite being much more versatile comparing with existing methods, our universal model still demonstrates comparable performance with the state-of-the-art finetuned methods. To the best of our knowledge, this is the first work that achieves modality-agnostic retinal vessel segmentation and also the first work that studies retinal vessel segmentation in some novel modalities.  [^1] Our codes and model will be available at the time of publication. We are also actively looking for approaches to release the three new vessel segmentation datasets. However, since retinal images and retinal vessels contain sensitive personal information, we can not make promise at this moment.
Index Terms:
retina, retinal image processing, vessel segmentation, image translation, topological learning.
@@ I Background and Introduction
Retinal vessel segmentation is a fundamental task in retinal image processing. It has multiple useful applications, including retinal image registration [1, 2], human biometric identification [3], retinal artery-vein classification [4] and tree topology analysis in digital images [5]. Clinically, the evaluation of retinal vessels is also important since the vessel features are predictive in several systemic diseases [6, 7, 8] and vessel segmentation is the key towards the automation of the evaluation. However, most works in retinal vessel segmentation focus on only the Color Fundus (CF) images. CF uses white visible light to illuminate the retina and the reflected light is used for generating the image. Nevertheless, although CF was the most widely-used modality in the past, it requires dilation before imaging and is uncomfortable for the patients. Additionally, the pathological information provided by CF is limited. Therefore, in more recent years, CF is used less frequently and emerging new imaging techniques become increasingly popular. These modalities include (Fig. 1): (1) Multi-Color Scanning Laser Ophthalmoscopy (MC), which uses more intense laser beam and allow exposing structures in deeper layers of the retina; (2) Florescence Angiography (FA), which uses florescence dye to illuminate the retinal vessels and are commonly used in the diagnosis of diseases related to neovascularization; (3) Fundus AutoFlorescence (FAF), which does not rely on external light source but uses natural fluorescence of the retina and provides insights into the health and metabolism of the retinal pigment epithelium (RPE) and photoreceptor cells; (4) Infrared/Near-Infrared Reflectance (IR/NIR), which uses Infrared or Near-Infrared lights to illuminate the image and allows seeing deeper features on the retina such as RPE abnormalities. (5) Optical Coherence Tomography Angiography (OCTA), which uses low-coherence interferometry to measure the backscattered light from retinal layers. OCTA has a much smaller imaging area than the other modalities and has considerably different retinal vessel morphology and topology.
@@
Figure 1: Overview of multi-modality retinal images and the segmentation results from our UVSM. In each image, the upper half is the original image and the lower half is the segmentation. In the middle, the camera (left) and modality (right) of the image is labeled. ‘Heidel’ refers to Heidelberg Spectralis.
@@
One major challenge that impedes the development of machine learning-based retinal vessel segmentation algorithms is the lack of data. Due to the complexity of the retinal vessels and the existence of many fine details, labeling is very time-consuming. Existing retinal vessel segmentation datasets (especially considering high label quality) are scarce and small even for the CF modality, which usually only have 20-40 images. In addition, to the best of our knowledge, there are no publicly available vessel segmentation datasets for FAF and IR. For FA, there is only a very small public dataset with 8 images [9]. In this work, we prepare three high-quality datasets for FA, FAF and IR, respectively. Each dataset has 40 images and the vessels are carefully labeled by retina experts. Nevertheless, despite the availability of the new datasets, in this work, we only use them for evaluation purpose. The reason is that: we believe that there can never be enough annotated images to allow training a powerful universal model. Instead, we need to train the model with limited existing data and seek to overcome the challenge through innovations in the algorithm. In fact, our segmentation network is trained with only 19 annotated CF images from the widely-used DRIVE [10] dataset.
@@
There are limited previous works which are extended beyond CF modality [11, 12, 13]. In addition to lacking comprehensive modalities (compared to our approach), one still needs to finetune a separate model on the target modality where each model only works for one image modality. We instead aim to build a single universal model that achieves robust segmentation for all the commonly-used modalities. The model could serve as an important tool in cross-modality retinal image study, which becomes increasingly important in modern research of the retina. To achieve the universal segmentation with limited labeled data, we develop our model in two major steps (Fig. 2). Firstly, we propose to use image translation as a preliminary domain adaptation method and train a model that translate input images from arbitrary modalities and cameras into a uniform Topcon (The Topcon TRC-50DX series is explicitly referred to as the gold standard in color fundus photography) CF image modality. We accessed abundant of de-identified multi-modality retinal images from multiple widely-used cameras (Topcon, Heidelberg Spectralis, Optos) at the Jacobs Retina Center at Shiley Eye Institute, University of California, San Diego (UCSD). Those images allow us to train a useful image translation model in a self-supervised approach. Furthermore, we adopt data augmentation to account for the slight domain difference with the other CF cameras which our center does not have, including Canon and Zeiss. Secondly, we train a segmentation model on the Topcon CF domain to perform the segmentation. However, image translation cannot fully eliminate the domain gaps between different modalities given such difficult many-to-one translation task. To solve this problem, we propose to use a topology-aware segmentation method that learns topological features of the vessels in addition to their conventional pixel-level features. The topology-aware segmentation methods were initially proposed to improve the topological accuracy of the segmentation of curvilinear structures. We go beyond their conventional applications and propose the topological learning as a novel domain adaptation method for retinal vessel segmentation, in a way that forces the segmentation to be topologically correct so when images from different domains are inputted, the model can still output robust segmentation.
@@
Our universal vessel segmentation model is evaluated on 10 diverse datasets (including 3 new datasets we prepared for this work) of the 6 most commonly-used retinal image modalities (CF, FA, FAF, MC, IR/NIR, OCTA) taken from 6 most commonly-used camera systems (Topcon, Canon, Zeiss, Heidelberg Spectralis, Optos, Optovue). Our universal model is compared with state-of-the-art methods which are finetuned on each of the 7 datasets, respectively, and is shown to achieve comparable performance. Extensive ablation studies were also performed to investigate how different components in the proposed pipeline affects the performance of the universal vessel segmentation. The main contributions of this paper are:
@@
  • We propose the universal retinal vessel segmentation task. To the best of our knowledge, this is the first work that treat retinal vessel segmentation as a modality-agnostic segmentation task covering multiple imaging modalities and camera systems, reflecting the real-world need for generalizable retinal vessel segmentation solutions.
@@
  • We propose the first universal (retinal) vessel segmentation model (UVSM) for multi-modality retinal images based on image translation and image topological learning.
@@
  • We investigate image-to-image translation between retinal images from different modalities. We explore different Generative Adversarial Nets (GAN)-based and Diffusion Model-based image translation methods and evaluate their performance on retinal images.
@@
  • We propose topological feature learning as a novel domain adaptation strategy for multi-modal retinal vessel segmentation. We show that by learning the topological features of the vessels, the segmentation model becomes much more robust to inputs from different domains (modalities and imaging systems).
@@
  • We collect and annotate three high-quality vessel segmentation datasets for the FA, FAF and IR retinal images, respectively.
@@
  • We provide a refined version of the widely-used DRIVE dataset with improved topological accuracy in the annotation (to be discussed in section IV-A2).
@@ II-A Retinal Vessel Segmentation
Retinal vessel segmentation is a specialized task that general-purpose foundational segmentation models [14, 15, 16] struggle to handle effectively. Current works in retinal vessel segmentation focus on segmenting vessels in Color Fundus retinal images, where various machine learning-based or non-machine learning-based methods were proposed. A majority of the early methods are based on the matched filter [17, 18, 19], which convolve retinal images with predefined or learned kernels and then adaptively threshold the image to obtain the vessel segmentation map. Another type of methods uses tracking algorithms to mathematically estimate the growth of the retinal vessel trees from seed points [20, 21]. The segmentation can then be derived by extending around the estimated vessel centerlines using the estimated vessel diameters. Some methods also adopt morphological image processing methods such as the top-hat operation to extract the vessel or as a post-processing step [22, 23]. Furthermore, numerous traditional machine learning-based methods were also applied in retinal vessel segmentation such as the support vector machine [24], random forest [25] and Adaboost [26] for supervised methods, and Gaussian mixture model [27], Fuzzy C-Means [28] for unsupervised methods. In more recent years, deep learning-based methods dominate the literature. They mainly rely on the Convolutional Neural Networks (CNN) as a backbone model due to the existence of vast fine details and the spread-out characteristic of the vessels. An introductory work is the Deep Retinal Image Understanding (DRIU) [29], which adopts a VGG [30] network structure as encoder and assigns specialized output layers to the embedded features to segment the vessels and the optic disc from the retina. Based on a widely-used U-Net [31] network structure and to seek further improvement for retinal vessel segmentation, [32, 33, 34, 35, 36] proposed to add various global reasoning modules at the bottom of the network. [35, 36, 37, 38, 34] proposed to add feature aggregation modules in the residual connection at each level of the network. [39, 40] proposed to use deformable convolution where the convolution kernel shapes are learned to adapt to the elongated morphology of the retinal vessels. With similar motivation, [41] proposed an orientation-selective convolution to learn selective receptive fields of the kernel to adapt to the shape of the vessels. [34] proposed to use improved feature fusion methods to better preserve the fine details of vessels in the segmentation. In addition to a single U-Net-like segmentation network, several works [42, 41, 43] proposed to use cascaded networks or a single looped network to iteratively refine the segmentation.
@@ II-B Unpaired Image-to-Image Translation
Image-to-image translation aims to transfer a source image to a target image domain while preserving its structural information. However, it is difficult to prepare paired retinal images from the source and target domain for training. Therefore, in this work we look into the unpaired image-to-image translation methods where the projection between the two domains can be learned with unpaired images with no structural correspondences. A representative and most-commonly used method in this task is CycleGAN [44], which introduces a cycle consistency that projects the translated image back to the source domain to force retaining the structural information in the translated image. Following works further introduce geometric consistency [45], mutual information regularization [46] and contrastive unpaired translation [47] to improve the computational efficiency over the cycle consistency-based methods during the training phrase. [48, 49] proposed one-to-many transfer schemes which allow translating images to multiple domains. While the above works are based on GAN backbone and achieved good image translation performance, diffusion models [50, 51, 52] have recently gained more research attention in this topic for easier training and superior image quality. However, the methods [53, 54, 55] are computationally intensive and were only applied on small images. For example, [55] combines diffusion model with adversarial learning and formulate the training in a cycle-consistency framework, which requires training two diffusion models at the same time. Most recently, [56] proposed an Unpaired Neuron Schrodinger Bridge (UNSB) with a novel scheme to simulate the Schrodinger Bridge, which is a random process that interpolates between two image domains, via a diffusion process. It has higher computational efficiency and makes application to higher-resolution images (e.g., retinal images) possible.
Image style transfer [57, 58, 59, 60, 61] is another method which can be seen as a special type of unpaired image-to-image translation but with a perceptual inductive bias. The majority of the methods use a perceptual feature space to compute a style loss and inject the style of a reference image to the input content image, while using another self-comparison content loss to retain the content information of the input in the transferred image. We also explore methods in this topic since style transfer is used for non-CF retinal vessel segmentation in [12, 13].
@@ II-C Topology-Aware Image Segmentation
Topology-aware image segmentation aims to preserve the topological accuracy in the segmentation of tubular and net-like structures. The literature mainly considers a topological loss function that extracts topological features from the ground truth and segmentation and minimizes the difference between them. One type of methods extracts certain features from the segmentation which are proximity to the actual topological features of the foreground objects. A pioneer work [62] uses an ImageNet-pretrained VGG [63, 30] network after the segmentation network. The perceptual features extracted from different layers of the pretrained VGG network is shown to be approximations of the topological features. The topological loss is then computed between the ground truth and prediction perceptual features. Another topological segmentation loss function [64] was proposed exclusively for the segmentation of glands. The loss is computed according to the iterations needed to derive the gland skeletons by the image erosion operation. [65] proposed to approximate the topological features of tubular structures by a relationship between the foregrounds in the ground truth and prediction and their soft-skeletons. Using such a relationship as a loss function is shown to strengthen the connectivity along the tubular structure thus providing better topological accuracy.
@@
Another type of methods uses strict topological features of the segmentation which are based on the computational topology theory and received more research attentions in recent years. The topological features are mainly based on the persistent homology theory [66, 67, 68, 69, 40, 70, 71, 72]. A revolutionary work that first introduces this concept in image segmenation is [66], which computes the persistent features from the ground truth and prediction, matching the features using a Wasserstein distance [73] and using the distance as the loss function. The following works share similar idea. [67] also uses the Wasserstein distance but apply the method on 3D point cloud data. [70] proposed to match the longest persistent barcodes in the prediction to the ground truth. [40] proposed to use a Hausdorff distance [74] to match the persistent diagrams to improve the sensitivity of the matching to the outliers. [72] and [71] both aim to compute more precise persistent feature matching, as in earlier methods the actual matching results are highly inaccurate due to only considering the global topological features. [72] and [71] instead allow the loss to act more selectively on the topological features thus improve the efficiency of the topological loss function. [71] proposed to use an Induced Matching [75] method which constructs a common ambient space between the ground truth and prediction and maps the persistent features in between to find a better matching. However, the computational cost of this method is very high, especially for large images, where typical retinal images usually have high resolutions. Alternatively, [72] proposed to leverage the spatial locations of the persistent features referencing their critical cell locations for more faithful persistent feature matching and is computationally much more efficient than [71].
@@
Figure 2: Pipeline for our proposed universal vessel segmentation model, which includes the 2-stage training and the inference.

### v4

@@ Abstract
We identify two major limitations in the existing studies on retinal vessel segmentation: (1) Most existing works are restricted to one modality, i.e., the Color Fundus (CF). However, multi-modality retinal images are used every day in the study of the retina and diagnosis of retinal diseases, and the study of vessel segmentation on other modalities is scarce; (2) Even though a few works extended their experiments to new modalities such as the Multi-Color Scanning Laser Ophthalmoscopy (MC), these works still require fine-tuning a separate model for the new modality. The fine-tuning will require extra training data, which is difficult to acquire. In this work, we present a novel universal vessel segmentation model (URVSM) for multi-modality retinal images. In addition to performing the study on a much wider range of image modalities, we also propose a universal model to segment the vessels in all these commonly used modalities. While being much more versatile compared with existing methods, our universal model also demonstrates comparable performance to the state-of-the-art fine-tuned methods. To the best of our knowledge, this is the first work that achieves modality-agnostic retinal vessel segmentation and the first to study retinal vessel segmentation in several novel modalities.  [^1] Code, model and 3 new retinal vessel segmentation datasets are available at https://github.com/JRC-VPLab/URVSM [^2] This paper is accepted by IEEE Transactions on Image Processing and will be published soon.
@@ I Background and Introduction
Retinal vessel segmentation is a fundamental task in retinal image processing. It has numerous important applications, including retinal image registration [1, 2], human biometric identification [3], retinal artery-vein classification [4] and tree topology analysis in digital images [5]. Clinically, the evaluation of retinal vessels is also important since the vessel features are predictive in several systemic diseases [6, 7, 8] and vessel segmentation is the key to the automation of the evaluation. However, most works in retinal vessel segmentation focus only on Color Fundus (CF) images. CF uses visible white light to illuminate the retina and the reflected light is used for generating the image. Nevertheless, although CF was the most widely used modality in the past, it requires dilation before imaging and is uncomfortable for patients. Moreover, the pathological information provided by CF is limited. Therefore, in more recent years, CF is used less frequently and emerging new imaging techniques become increasingly popular. These modalities include (Fig. 1): (1) Multi-Color Scanning Laser Ophthalmoscopy (MC), which uses a more intense laser beam and allows visualization of structures in deeper layers of the retina; (2) Fluorescence Angiography (FA), which uses fluorescent dye to illuminate the retinal vessels and is commonly used in the diagnosis of diseases related to neovascularization; (3) Fundus AutoFluorescence (FAF), which does not rely on an external light source but uses natural fluorescence of the retina and provides insights into the health and metabolism of the retinal pigment epithelium (RPE) and photoreceptor cells; (4) Infrared/Near-Infrared Reflectance (IR/NIR), which uses Infrared or Near-Infrared light to illuminate the image and allows visualization of deeper features on the retina, such as RPE abnormalities; (5) Optical Coherence Tomography Angiography (OCTA), which uses low-coherence interferometry to measure the backscattered light from retinal layers. OCTA has a much smaller imaging area than the other modalities and has a considerably different retinal vessel morphology and topology.
@@
Figure 1: Overview of multi-modality retinal images and the segmentation results from our URVSM. In each image, the upper half is the original image and the lower half is the segmentation. In the middle, the camera (left) and modality (right) of the image is labeled. ‘Heidel’ refers to Heidelberg Spectralis.
@@
One major challenge that impedes the development of machine learning-based retinal vessel segmentation algorithms is the lack of data. Due to the complexity of the retinal vessels and the existence of many fine details, labeling is very time-consuming. Existing retinal vessel segmentation datasets (especially those with high label quality) are scarce and small, even for the CF modality, which usually only has 20-40 images. In addition, to the best of our knowledge, there are no publicly available vessel segmentation datasets for FAF and IR. For FA, there is only a very small public dataset with 8 images [9]. In this work, we prepare three high-quality datasets for FA, FAF and IR, respectively. Each dataset has 40 images, and the vessels are carefully labeled by retinal experts. Nevertheless, despite the availability of the new datasets, in this work, we use them only for evaluation purposes. The reason is that we believe that there can never be enough annotated images to allow the training of a powerful universal model. Instead, we need to train the model with limited existing data and seek to overcome the challenge through innovations in the algorithm. In fact, our segmentation network is trained with only 19 annotated CF images from the widely used DRIVE [10] dataset.
@@
There are limited previous works that extend beyond the CF modality [11, 12, 13]. In addition to lacking comprehensive modalities (compared to our approach), one still needs to fine-tune a separate model on the target modality, where each model only works for one image modality. We instead aim to build a single universal model that achieves robust segmentation for all the commonly used modalities. The model could serve as an important tool in the cross-modality retinal image research, which becomes increasingly important in modern research of the retina. To achieve the universal segmentation with limited labeled data, we develop our model in two major steps (Fig. 2). First, we propose to use image translation as a preliminary domain adaptation method and train a model that translates input images from arbitrary modalities and cameras into a uniform Topcon (the Topcon TRC-50DX series is explicitly referred to as the gold standard in color fundus photography) CF image modality. We accessed an abundance of de-identified multi-modality retinal images from multiple widely used cameras (Topcon, Heidelberg Spectralis, Optos) at the Jacobs Retina Center at Shiley Eye Institute, University of California, San Diego (UCSD). These images allow us to train a useful image translation model in a self-supervised manner. Furthermore, we adopt data augmentation to account for the slight domain difference with the other CF cameras which our center does not have, including Canon and Zeiss. Secondly, we train a segmentation model on the Topcon CF domain to perform the segmentation. However, image translation cannot fully eliminate the domain gaps between different modalities given such a difficult many-to-one translation task. To solve this problem, we propose to use a topology-aware segmentation method that learns topological features of the vessels in addition to their conventional pixel-level features. The topology-aware segmentation methods were initially proposed to improve the topological accuracy of the segmentation of curvilinear structures. We go beyond their conventional applications and propose topological learning as a novel domain adaptation method for retinal vessel segmentation, in a way that forces the segmentation to be topologically correct, so when images from different domains are input, the model can still yield robust segmentation.
@@
Our universal vessel segmentation model is evaluated on 10 diverse datasets (including 3 new datasets we prepared for this work) of the 6 most commonly-used retinal image modalities (CF, FA, FAF, MC, IR/NIR, OCTA) taken from 6 most commonly used camera systems (Topcon, Canon, Zeiss, Heidelberg Spectralis, Optos, Optovue). Our universal model is compared with state-of-the-art methods that are fine-tuned on each of the 7 datasets, respectively, and is shown to achieve comparable performance. Extensive ablation studies were also performed to investigate how different components in the proposed pipeline affect the performance of the universal vessel segmentation. The main contributions of this paper are:
@@
  • We propose the universal retinal vessel segmentation task. To the best of our knowledge, this is the first work that treats retinal vessel segmentation as a modality-agnostic segmentation task covering multiple imaging modalities and camera systems, reflecting the real-world need for generalizable retinal vessel segmentation solutions.
@@
  • We propose the first universal (retinal) vessel segmentation model (URVSM) for multi-modality retinal images.
@@
  • We investigate image-to-image translation between retinal images from different modalities. We explore different Generative Adversarial Network (GAN)-based and Diffusion Model-based image translation methods and evaluate their performance on retinal images.
@@
  • We propose topological feature learning as a novel domain adaptation strategy for multi-modal retinal vessel segmentation. We show that by learning a combination of established topological features of the vessels, the segmentation model becomes much more robust to inputs from different domains (modalities and imaging systems).
@@
  • We release three new, high-quality vessel segmentation datasets for the FA, FAF, and IR retinal images, respectively.
@@
  • We release refined vessel annotations of the widely used DRIVE dataset with improved topological accuracy (to be discussed in section IV-A2).
@@ II-A Retinal Vessel Segmentation
Retinal vessel segmentation is a specialized task that general-purpose foundational segmentation models [14, 15, 16] struggle to handle effectively. Current works in retinal vessel segmentation focus on segmenting vessels in Color Fundus retinal images, where various machine learning-based or non-machine learning-based methods were proposed. A majority of the early methods are based on match filtering [17, 18, 19], which convolves retinal images with predefined or learned kernels and then adaptively thresholds the image to obtain the vessel segmentation map. Another type of methods use tracking algorithms to mathematically estimate the growth of the retinal vessel trees from seed points [20, 21]. The segmentation can then be derived by extending around the estimated vessel centerlines using the estimated vessel diameters. Some methods also adopt morphological image processing methods such as the top-hat operation to extract the vessels or as a post-processing step [22, 23]. Furthermore, numerous traditional machine learning-based methods were also applied in retinal vessel segmentation such as the support vector machine [24], random forest [25] and Adaboost [26] for supervised methods, and the Gaussian mixture model [27], Fuzzy C-Means [28] for unsupervised methods. In more recent years, deep learning-based methods dominate the literature. They mainly rely on the Convolutional Neural Networks (CNN) as a backbone models due to the existence of vast fine details and the spread-out characteristics of the vessels. An introductory work is the Deep Retinal Image Understanding (DRIU) [29], which adopts a VGG [30] network structure as an encoder and assigns specialized output layers to the embedded features to segment the vessels and the optic disc from the retina. Based on a widely used U-Net [31] network structure and to seek further improvement for retinal vessel segmentation, [32, 33, 34, 35, 36] proposed to add various global reasoning modules at the bottom of the network. [35, 36, 37, 38, 34] proposed to add feature aggregation modules in the residual connection at each level of the network. [39, 40] proposed to use deformable convolution where the convolution kernel shapes are learned to adapt to the elongated morphology of the retinal vessels. With similar motivation, [41] proposed an orientation-selective convolution to learn selective receptive fields of the kernel to adapt to the shape of the vessels. [34] proposed to use improved feature fusion methods to better preserve the fine details of vessels in the segmentation. In addition to a single U-Net-like segmentation network, several works [42, 41, 43] proposed to use cascaded networks or a single looped network to iteratively refine the segmentation.
@@ II-B Unpaired Image-to-Image Translation
Image-to-image translation aims to transfer a source image to a target image domain while preserving its structural information. However, it is difficult to prepare paired retinal images from the source and target domains for training. Therefore, in this work, we look into the unpaired image-to-image translation methods where the projection between the two domains can be learned with unpaired images with no structural correspondences. A representative and most commonly used method in this task is CycleGAN [44], which introduces a cycle consistency that projects the translated image back to the source domain to enforce retaining the structural information in the translated image. Following works further introduce geometric consistency [45], mutual information regularization [46], and contrastive unpaired translation [47] to improve the computational efficiency over the cycle consistency-based methods during the training phase. [48, 49] proposed one-to-many transfer schemes which allow the translation of images to multiple domains. While the above works are based on GAN backbones and achieve good image translation performance, diffusion models [50, 51, 52] have recently gained more research attention on this topic for easier training and superior image quality. However, the methods [53, 54, 55] are computationally intensive and were only applied to small images. For example, [55] combines a diffusion model with adversarial learning and formulates the training in a cycle-consistency framework, which requires training two diffusion models at the same time. Most recently, [56] proposed an Unpaired Neuron Schrödinger Bridge (UNSB) with a novel scheme to simulate the Schrodinger Bridge, which is a random process that interpolates between two image domains via a diffusion process. It has higher computational efficiency and makes the application to higher-resolution images (e.g., retinal images) possible.
Image style transfer [57, 58, 59, 60, 61] is another method which can be seen as a special type of unpaired image-to-image translation but with a perceptual inductive bias. The majority of the methods use a perceptual feature space to compute a style loss and inject the style of a reference image into the input content image, while using another self-comparison content loss to retain the content information of the input in the transferred image. We also explore methods in this topic since style transfer is used for non-CF retinal vessel segmentation in [12, 13].
@@ II-C Topology-Aware Image Segmentation
Topology-aware image segmentation aims to preserve the topological accuracy in the segmentation of tubular and net-like structures. The literature mainly considers a topological loss function that extracts topological features from the ground truth and segmentation and minimizes the difference between them. One type of methods extract certain features from the segmentation which approximate the actual topological features of the foreground objects. A pioneering work [62] uses an ImageNet-pretrained VGG [63, 30] network after the segmentation network. The perceptual features extracted from different layers of the pretrained VGG network are shown to be approximations of the topological features. The topological loss is then computed between the ground truth and predicted perceptual features. Another topological segmentation loss function [64] was proposed exclusively for the segmentation of glands. The loss is computed according to the iterations needed to derive the gland skeletons by the image erosion operation. [65] proposed to approximate the topological features of tubular structures by a relationship between the foregrounds in the ground truth and prediction and their soft-skeletons. Using such a relationship as a loss function is shown to strengthen the connectivity along the tubular structure, thus providing better topological accuracy.
@@
Another type of methods use strict topological features of the segmentation which are based on the computational topology theory and have received more research attention in recent years. The topological features are mainly based on the persistent homology theory [66, 67, 68, 69, 40, 70, 71, 72]. A revolutionary work that first introduced this concept in image segmentation is [66], which computes the persistent features from the ground truth and prediction, matching the features using a Wasserstein distance [73] and using the distance as the loss function. The following works share a similar idea. [67] also uses the Wasserstein distance but applies the method on 3D point cloud data. [70] proposed to match the longest persistent barcodes in the prediction to the ground truth. [40] proposed to use a Hausdorff distance [74] to match the persistent diagrams to improve the sensitivity of the matching to the outliers. [72] and [71] both aim to compute more precise persistent feature matching, as in earlier methods, the actual matching results are highly inaccurate due to only considering the global topological features. [72] and [71] instead allow the loss to act more selectively on the topological features thus improves the efficiency of the topological loss function. [71] proposed to use an Induced Matching [75] method which constructs a common ambient space between the ground truth and prediction and maps the persistent features in between to find a better matching. However, the computational cost of this method is very high, especially for large images, where typical retinal images usually have high resolutions. Alternatively, [72] proposes to leverage the spatial locations of the persistent features referencing their critical cell locations for more faithful persistent feature matching, and is computationally much more efficient than [71].
@@
Figure 2: Pipeline for our proposed universal retinal vessel segmentation model, which includes the 2-stage training and the inference.